Also, what is the NPV? using the cash flowA store will cost $625,000 to open. Variable costs will be 38% of sales and fixed costs are $250,000 per year. The investment
Answer #1

700,000 =700000*38%+250000 =625000/9 Sales Less: Costs Depreciation EBIT Less: Tax @ 35% Net income Add: Depreciation 516,000

*Please rate thumbs up